OpenStudy (anonymous):

Solve 2/5x = -3/15 + 1/x

OpenStudy (anonymous):

help anyone????

OpenStudy (kirbykirby):

Just to be clear, is the first term \[\frac{2}{5x}\text{ or }\frac{2}{5}x\]

OpenStudy (anonymous):

2/5x first one right

OpenStudy (anonymous):

first you need to get all x terms in one side so you can solve for it

OpenStudy (anonymous):

\[\frac{ 2 }{ 5x } - \frac{ 1 }{ x } = \frac{ -3 }{ 15 }\] can you do it from there?

OpenStudy (anonymous):

whats the steps

OpenStudy (anonymous):

you need to find a common denominator so you can subtract 2/5x - 1/x

OpenStudy (kirbykirby):

\[ \frac{2}{5x}-\frac{1}{x}=-\frac{3}{15}\\ \, \\ \frac{2}{5x}-\frac{5}{5x}=-\frac{3}{15} \\ \, \\ \frac{2-5}{5x}=-\frac{3}{15}\\ \, \\ \frac{-3}{5x}=-\frac{3}{15}\\ \, \\ -3(15)=-3(5x)\\ \,\\ 45=15x \\ \, \\ x=3\]

OpenStudy (anonymous):

thanks